Find parametric equations through point P= (2,2,7) in the direction of the vector v = (44, 14, -20).
Answer by ikleyn(52781)

.
In 3D space these parametric equations of a straight line are

    x - 2 = 44t,

    y - 2 = 14t,

    z - 7 = -20t,


where t is the parameter.